# A Framework for Lorentz-Dirac Dynamics and Post-Newtonian Interaction of Radiating Point Charges

### v4

@@ Appendix C Details on the Darwin force correction at 1PN
In this section, we show the detailed steps leading to Eq. (63) of the main text. We start from the Darwin interaction Hamiltonian, along with the useful notations as presented below:
@@
The velocity–dependent force derived in Eq. (63) of the main text is the classical, spin–independent component of the Darwin interaction. It represents the leading relativistic correction to Coulomb’s law obtained by expanding the full electromagnetic interaction between two charges to order $v^{2}/c^{2}$. This correction encodes conservative retardation effects in the near zone and admits three equivalent interpretations.
@@
Appendix D Closer look into the Four–acceleration and the Covariant Larmor Power
In this appendix we derive the explicit form of the four–acceleration $a^{\mu}=du^{\mu}/d\tau$ in terms of the ordinary velocity $\mathbf{v}$ and acceleration $\mathbf{a}=d\mathbf{v}/dt$, and we show how the covariant expression for the radiated power reduces to the standard Liénard generalization of the Larmor formula. The four–velocity of a particle with ordinary velocity $\mathbf{v}$ is
$u^{\mu}=\gamma(c,\mathbf{v}),\qquad\gamma=\frac{1}{\sqrt{1-v^{2}/c^{2}}}.$ (113)
Since $d\tau=dt/\gamma$, the derivative with respect to proper time may be written as
$\frac{d}{d\tau}=\gamma\frac{d}{dt}.$ (114)
We first compute the time component. Because $u^{0}=\gamma c$, we obtain
$a^{0}=\frac{du^{0}}{d\tau}=\gamma\frac{d}{dt}(\gamma c)=\gamma c\frac{d\gamma}{dt}.$ (115)
Using
$\frac{d\gamma}{dt}=\gamma^{3}\frac{\mathbf{v}\cdot\mathbf{a}}{c^{2}},$ (116)
we find
$a^{0}=\gamma^{4}\frac{\mathbf{v}\cdot\mathbf{a}}{c}.$ (117)
For the spatial components, begin with $\mathbf{u}=\gamma\mathbf{v}$. Then
$\mathbf{a}^{(4)}$ $=\frac{d\mathbf{u}}{d\tau}=\gamma\frac{d}{dt}(\gamma\mathbf{v})=\gamma(\gamma\mathbf{a}+\mathbf{v}\,\dot{\gamma}),$ (118)
$\dot{\gamma}$ $=\gamma^{3}\frac{\mathbf{v}\cdot\mathbf{a}}{c^{2}}.$
This gives
$\mathbf{a}^{(4)}=\gamma^{2}\mathbf{a}+\gamma^{4}\frac{\mathbf{v}\cdot\mathbf{a}}{c^{2}}\mathbf{v}.$ (119)
Combining Eqs. (117) and (119), the four–acceleration takes the explicit form
$a^{\mu}=\left(\gamma^{4}\frac{\mathbf{v}\cdot\mathbf{a}}{c},\;\gamma^{2}\mathbf{a}+\gamma^{4}\frac{\mathbf{v}\cdot\mathbf{a}}{c^{2}}\mathbf{v}\right).$ (120)
Using the metric signature (,+,+,+)(-,+,+,+),
$a_{\mu}a^{\mu}=-(a^{0})^{2}+|\mathbf{a}^{(4)}|^{2}.$ (121)
Substituting Eq. (120) and simplifying, one obtains the standard invariant form
$a_{\mu}a^{\mu}=\gamma^{6}\!\left(a^{2}-\frac{(\mathbf{v}\times\mathbf{a})^{2}}{c^{2}}\right)=\gamma^{4}\!\left(a^{2}+\gamma^{2}\frac{(\mathbf{v}\cdot\mathbf{a})^{2}}{c^{2}}\right).$ (122)
Both expressions are equivalent via $(\mathbf{v}\times\mathbf{a})^{2}=v^{2}a^{2}-(\mathbf{v}\cdot\mathbf{a})^{2}$. The covariant form of the radiated power for a particle with rest mass $m$ and charge $q$ is
$P=-\frac{2}{3}\frac{q^{2}}{4\pi\epsilon_{0}m^{2}c^{3}}\left(\frac{dp_{\mu}}{d\tau}\frac{dp^{\mu}}{d\tau}\right),\qquad p^{\mu}=mu^{\mu}.$ (123)
Since $dp^{\mu}/d\tau=ma^{\mu}$, the factors of $m$ cancel and we obtain
$P=-\frac{2}{3}\frac{q^{2}}{4\pi\epsilon_{0}c^{3}}\,a_{\mu}a^{\mu}.$ (124)
Substituting Eq. (122), the radiated power becomes
$P=\frac{q^{2}}{6\pi\epsilon_{0}c^{3}}\,\gamma^{6}\!\left(a^{2}-\frac{(\mathbf{v}\times\mathbf{a})^{2}}{c^{2}}\right)\\$ (125)
or equivalently,
$P=\frac{q^{2}}{6\pi\epsilon_{0}c^{3}}\,\gamma^{4}\!\left(a^{2}+\gamma^{2}\frac{(\mathbf{v}\cdot\mathbf{a})^{2}}{c^{2}}\right).$ (126)
In the nonrelativistic limit $\gamma\to 1$ and $v\ll c$, both forms reduce to the familiar Larmor result,
$P=\frac{q^{2}a^{2}}{6\pi\epsilon_{0}c^{3}}.$ (127)
Appendix E Exact Solution of the Landau–Lifshitz Equation in a Coulomb Potential
In this appendix we summarize the exact solution of the non–relativistic Landau–Lifshitz (LL) equation for a charged particle moving in a Coulomb potential, following the analysis of Rajeev [37]. This provides a rare example of a dissipative classical system that remains integrable and admits an exact analytic solution, expressed in terms of Painlevé solutions. The Landau–Lifshitz equation of motion for a radiating charged particle in an electrostatic potential $\Phi(\mathbf{r})$ reduces, in the non–relativistic limit, to
$\frac{d}{dt}\!\left(\mathbf{v}+s\,\nabla U\right)+\nabla U=0,$ (128)
where
$U=\frac{q}{m}\Phi,\qquad s=\frac{2}{3}\frac{q^{2}}{mc^{3}}$ (129)
is the characteristic radiation–reaction time scale. Equation (128) is second order in time and avoids the runaway solutions of the Lorentz–Dirac equation. For a central potential $U(r)$,
$\nabla U=\hat{\mathbf{r}}\,U_{r},\qquad U_{r}\equiv\frac{dU}{dr},$ (130)
and the equation of motion becomes
$\frac{d\mathbf{v}}{dt}+s\,\frac{d}{dt}\!\left(\hat{\mathbf{r}}U_{r}\right)+\hat{\mathbf{r}}U_{r}=0.$ (131)
Defining the angular momentum $\mathbf{L}=\mathbf{r}\times\mathbf{v}$, we find
$\frac{d\mathbf{L}}{dt}=-s\,\frac{U_{r}}{r}\,\mathbf{L}.$ (132)
Thus the direction of $\mathbf{L}$ is conserved, while its magnitude decays monotonically. The motion is therefore planar, and for nonzero initial angular momentum the trajectory spirals inward. For a Coulomb potential
$U(r)=\frac{k}{r},\qquad U_{r}=-\frac{k}{r^{2}},$ (133)
Eq. (132) becomes
$\frac{dL}{dt}=\frac{ks}{r^{3}}L.$ (134)
A key simplification arises from the identity
$\frac{L^{2}}{r^{3}}=\frac{1}{2ks}\frac{d}{dt}\!\left(L^{2}\right),$ (135)
which allows the radial equation to be written as a total derivative.
Introducing the radial velocity $v_{r}=\dot{r}$ and defining
$z\equiv v_{r}-\frac{ks}{r^{2}}-\frac{L^{2}}{2ks},$ (136)
the equations of motion reduce to an autonomous system,
$\frac{dL}{dt}$ $=\frac{ks}{r^{3}}L,$
$\frac{dz}{dt}$ $=\frac{k}{r^{2}},$
$\frac{dr}{dt}$ $=\frac{L^{2}}{2ks}+z+\frac{ks}{r^{2}}.$
Eliminating the time variable in favor of $z$, one obtains a closed second order equation for the angular momentum:
$\frac{d^{2}L}{dz^{2}}=-\frac{1}{2k^{2}}L^{3}-\frac{s}{k}zL.$ (137)
Up to rescalings, Eq. (137) is the Painlevé II equation (with parameter $a=0$),
$\frac{d^{2}u}{dx^{2}}=2u^{3}+xu.$ (138)
Defining
$b=\left(-\frac{s}{k}\right)^{1/3},\qquad a=\sqrt{-2k^{2}}\,b,$ (139)
the solution can be written compactly as
L(z)=au(bz),L(z)=a\,u(bz), (140)
where $u(x)$ is a Painlevé II transcendent. For an attractive Coulomb potential ($k<0$), the physically relevant solutions are purely imaginary $u(x)$ with real argument $x=bz$. Once $u(z)$ is known, the full trajectory follows from quadratures. Using
$\frac{dz}{dt}=\frac{k}{r^{2}},\qquad L=r^{2}\frac{d\theta}{dt},$ (141)
one finds
$r(z)$ $=s\,\frac{u(z)}{u^{\prime}(z)},$ (142)
$\theta(z)$ $=\theta_{0}+\frac{a}{k}\int^{z}u(bz^{\prime})\,dz^{\prime}.$ (143)
Thus the orbit is given parametrically by $\{r(z),\theta(z)\}$. The Painlevé II solutions describe both slowly decaying bound orbits and capture trajectories. In particular, for attractive Coulomb interaction, generic initial conditions lead to a monotonic decrease of LL and $r$, implying that the charged particle spirals inward and is captured in finite time. This confirms that the Landau–Lifshitz equation yields physically sensible radiation–driven inspiral behavior, in contrast to the runaway solutions of the Lorentz–Dirac equation. The appearance of Painlevé transcendents indicates a hidden integrable structure in this dissipative system, with conserved quantities replacing energy and angular momentum. This structure is closely analogous to radiation–reaction–driven inspirals in gravitational two–body dynamics and provides a useful analytic model for more complicated self–force problems [39, 40, 41].

### v5

@@ Appendix C Details on the Darwin force correction at 1PN
In this section, we show the detailed steps leading to the 1PN conservative force in Eq. (17) of the main text. We start from the Darwin interaction Hamiltonian, along with the useful notations as presented below:
@@
The velocity–dependent conservative 1PN force derived in Eq. (17) of the main text is the classical, spin–independent component of the Darwin interaction. It represents the leading relativistic correction to Coulomb’s law obtained by expanding the full electromagnetic interaction between two charges to order $v^{2}/c^{2}$. This correction encodes conservative retardation effects in the near zone and admits three equivalent interpretations.
@@
Appendix D Simulation results: the Charge Neutral Binary
D.1 The conservative sector
(a) Trajectories.
(b) Darwin Hamiltonian evolution.
(a) 3D trajectories with initial (green) and final (red) markers.
(b) Parametric projection in the xxyy plane.
(c) $x(\tau)$ vs. $\tau$.
(d) $y(\tau)$ vs. $\tau$.
(e) $z(\tau)$ vs. $\tau$.
(f) Hamiltonian $H(\tau)$ showing radiative energy loss.
Figure 5: Charge neutral binary of same mass starting from diametrically opposite locations with equal and opposite momenta, with radiation reaction from the truncated post Newtonian approach: (a) Full 3D trajectories; (b) parametric xxyy plane; (c–e) components $x(\tau)$, $y(\tau)$, $z(\tau)$; (f) Hamiltonian $H(\tau)$, decreasing due to radiation. The initial particle positions are shown in green, and the final positions at $\tau=\tau_{f}$ are shown in red.
(a) 3D trajectories with initial (green) and final (red) markers.
(b) Parametric projection in the xxyy plane.
(c) $x(\tau)$ vs. $\tau$.
(d) $y(\tau)$ vs. $\tau$.
(e) $z(\tau)$ vs. $\tau$.
(f) Hamiltonian $H(\tau)$ showing radiative energy loss.
Figure 6: Charge neutral binary of extreme mass ratio, one heavy and one light (with circular orbit initial conditions), with radiation reaction from the truncated post Newtonian approach: (a) Full 3D trajectories; (b) parametric xxyy plane; (c–e) components $x(\tau)$, $y(\tau)$, $z(\tau)$; (f) Hamiltonian $H(\tau)$, decreasing due to radiation. The initial particle positions are shown in green, and the final positions at $\tau=\tau_{f}$ are shown in red.
(a) 3D trajectories with initial (green) and final (red) markers.
(b) Parametric projection in the xxyy plane.
(c) Hamiltonian $H(\tau)$ showing radiative energy loss.
(d) $x(\tau)$ vs. $\tau$.
(e) $y(\tau)$ vs. $\tau$
(f) $\frac{\mathcal{I}}{\mathcal{I}_{0}}$ vs No. of orbits
Figure 7: Charge neutral binary, one heavy and one light, with elliptic orbit initial conditions. Radiation reaction from the truncated post Newtonian approach shows clear signature of circularization: (a) Full 3D trajectories; (b) parametric xxyy plane; (c) Hamiltonian $H(\tau)$, decreasing due to radiation (d-e) components $x(\tau)$, $y(\tau)$ and (f) plot showing $\mathcal{I}=a(1-e^{2})/e^{4/3}$ is preserved to sub-percent accuracy, with $\mathcal{O}(1/c^{2})$ corrections as expected . The initial particle positions are shown in green, and the final positions at $\tau=\tau_{f}$ are shown in red.
As a stringent validation of the conservative sector of our framework, we consider a system of two spatially separated three–dimensional charge blobs of opposite sign, evolved using the Darwin Hamiltonian alone, with all radiation–reaction terms explicitly switched off. Each blob consists of multiple point charges initialized within a compact spherical region, with the two charge clouds separated by a distance large compared to their individual sizes.
The corresponding Hamiltonian $H(\tau)$ is shown in Fig. 4(b). As expected for purely conservative dynamics, no secular drift is observed. The small bounded fluctuations visible in $H(\tau)$ arise from unavoidable numerical stiffness during close encounters between particles and provide a diagnostic of numerical accuracy. The absence of any systematic energy loss confirms that the Darwin sector has been implemented consistently and provides a reliable starting point for the inclusion of dissipative radiation–reaction effects in the next section. We now present representative solutions of the dissipative 1.5PN accurate NN–body framework Eq. (16). Throughout this section we use $\tau$ to denote the ordinary coordinate time (not the proper time). The conservative interactions are evolved with the Darwin Hamiltonian, retaining all terms through 1PN order, $\mathcal{O}(1/c^{2})$, while dissipation is incorporated as a non-Hamiltonian Landau–Lifshitz radiation–reaction force at leading 1.5PN order, $\mathcal{O}(1/c^{3})$, implemented via order reduction. This canonical-plus-dissipative framework allows us to test how radiative losses modify otherwise integrable Coulombic motion, producing secular inspiral and circularization. The examples below progress from a symmetric equal-mass neutral binary to an extreme-mass-ratio “hydrogen-like” configuration and finally to an initially eccentric orbit that circularizes under radiation reaction; in each case we monitor trajectories and the Hamiltonian $H(\tau)$ to quantify the cumulative energy loss.
D.2 Charge neutral binary of same mass
$\mathbf{r}_{1}(0)=(1,0,0),\qquad\mathbf{r}_{2}(0)=(-1,0,0),$
$\frac{mv^{2}}{R}=\frac{\alpha}{(2R)^{2}},$
$\mathbf{p}_{1}(0)=(0,\,0.5,\,0),\qquad\mathbf{p}_{2}(0)=(0,\,-0.5,\,0),$
$E_{\rm tot}=mv^{2}-\frac{k}{2R}=-\,\frac{1}{4},$
D.3 Charge neutral binary of extreme mass ratio
with momenta
D.4 Charge neutral binary of comparable mass showing circularization of initial elliptic orbits
In Fig. 1 we examine a two–body configuration demonstrating the radiative circularization predicted by the truncated post-Newtonian scheme Eq. (16). We consider oppositely charged particles with masses $(m_{1},m_{2})=(5,1)$ initialized in the center–of–mass frame at a separation $r_{0}=1$,